You have been asked by the king of the small island nation of Littleston to present the labor force data listed below.
 

Number of employedWorking-age populationLabor force
20,00040,00032,000 


Using this data, calculate the number of unemployed people, the unemployment rate, and the labor force participation rate.

Enter your answers in the table below. (Give all answers to two decimals.)

Number of unemployed peopleUnemployment rateLabor force participation rate
          %        %   
Answer #1

Number of unemployed people = Labor force - employed = 32000 - 20000 = 12000


Unemployed rate = unemployed / labor force * 100 = 12000 / 32000 * 100 = 37.50%


Labor force participation rate = Labor force / Working age * 100 = 32000 / 40000 * 100 = 80%


answered by: feng